OpenStudy (anonymous):

determine whether the following function is a linear transformation,justify your answer. T:V>R,where V is an inner product space and T(U)=//U//

OpenStudy (turingtest):

the transformation is\[T(\vec u)=\|\vec u\|\]?

OpenStudy (anonymous):

yep

OpenStudy (turingtest):

you know the two tests we need to do to see if this is a linear transformation?

OpenStudy (anonymous):

yes i do

OpenStudy (turingtest):

\[T(\vec u+\vec v)=T(\vec u)+T(\vec v)\]and\[T(c\vec u)=cT(\vec u)\]

OpenStudy (turingtest):

so let's try the second one first; seems easier....

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

\[\vec u=\langle a_1,a_2,...a_n\rangle\]\[c\vec u=\langle ca_1,ca_2,...ca_n\rangle\]\[T(c\vec u)=\sqrt{c^2a_1^2+c^2a_2^2+...+c^2a_n^2}\]you tell me, is that the same as\[cT((\vec u)\]?

OpenStudy (anonymous):

absolutely the same

OpenStudy (turingtest):

I agree, so on to the other test...

OpenStudy (turingtest):

\[\vec u+\vec v=\langle a_1+b_1,a_2+b_2,...,a_n+b_n\rangle\]\[T(\vec u+\vec v)=\sqrt{(a_1+b_1)^2+(a_2+b_2)^2+...+(a_n+b_n)^2}\]is that equal to\[\sqrt{a_1^2+a_2^2+...+a_n^2}+\sqrt{b_1^2+b_2^2+...+b_n^2}\]?

OpenStudy (anonymous):

in nt sure

OpenStudy (turingtest):

\[T(\vec u+\vec v)=\sqrt{(a_1+b_1)^2+(a_2+b_2)^2+...+(a_n+b_n)^2}\]\[=\sqrt{a_1^2+2a_1b_1+b_1^2+a_2^2+2a_2b_2+b_2^2+...+a_n^2+2a_nb_n+b_n^2}\]looks to me like we sonw be able to combine the radicals in\[\sqrt{a_1^2+a_2^2+...+a_n^2}+\sqrt{b_1^2+b_2^2+...+b_n^2}\]or get rid of those cross terms \(2a_kb_k\), so I would say they seem not to be the same

OpenStudy (phi):

use u= -v

OpenStudy (turingtest):

oh yeah that would have made it way more obvious :P

OpenStudy (anonymous):

so it doesnt satisfy the two conditions?

OpenStudy (turingtest):

try phi's suggestion and it should be clear that it does not satisfy the last condition

OpenStudy (turingtest):

\[\vec v=-\vec u\]\[T(\vec u+(-\vec u))=T(0)=0\neq\|\vec u\|+\|-\vec u\|=2\|\vec u\|\]hence\[T(\vec u+\vec v)\neq T(\vec u)+T(\vec v)\]

OpenStudy (anonymous):

meaning is not a linear transformation?

OpenStudy (turingtest):

if it fail either test it is not a linear transformation

OpenStudy (turingtest):

...so yeah